Suppose we are testing Ho: p=.20 vs Ha: p<.20 and TS=2.34. What is the p-value?

0.0096

0.9904

0.0107

0.0192

0.9893

0.0214
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is

H0 : p = 0.20

Ha : p <  0.20

Test statistic = z = 2.34

P(z > 2.34) = 1 - P(z < 2.34) = 1 - 0.9904

P-value = 0.0096